OpenStudy (anonymous):

a boat travels against a stream flowing at 3km/h for 20km and then return to its starting pt with the current.if the whole journey takes 7 hours,find the speed of the boat in still water

OpenStudy (anonymous):

t = d/r Let r be the speed of the boat in still water. Solve the following for r:\[\frac{20}{r-3}+\frac{20}{r+3}=7 \]r = 7 km/h